OpenStudy (anonymous):

The fifth term of the expansion (1-pi)^6 is...

OpenStudy (anonymous):

pi or \(\pi \) ?

OpenStudy (anonymous):

\[\dbinom{6}{1}(-1)\pi^5=-6\pi^5\]

OpenStudy (anonymous):

@satellite73: I think that is the 6th term.

OpenStudy (cwrw238):

formula worth remembering is: (r+1)th term of (a +x) ^n = nCr a^(n-r) x^r